gcarson:

What are the approximate solutions of 5x2 − 7x = 1, rounded to the nearest hundredth?

Tranquility:

Hello @gcarson Welcome to QuestionCove!

Tranquility:

Do you know how to solve quadratic equations?

Tranquility:

Bring the 1 to the other side of the equation so you have it in the form of ax^2 + bx + c = 0 And then use the quadratic formula \(\Large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)